OpenStudy (anonymous):

Write all two-factor factorizations of each number. Then list all the factors of the number. 1. 50 2. 42

OpenStudy (anonymous):

50 = 1*50 2*25 5*10 42 = 1*42 2*21 3*14 6*7

OpenStudy (anonymous):

Thank you :)